OpenStudy (christos):

Vectors A = 3i - 2j + 4k B = -5i + 2j - k A x 4B = -24i-68j-16k PLEASE VERIFY ?

OpenStudy (anonymous):

A=3i-2j+4k 4B=-20i+8j-4k i j k 3 -2 4 -20 8 -4 +8i-80j+24k-32i+12j-40k -24i-68j-16k Yep.

OpenStudy (christos):

Then why my teacher gets this ? http://screencast.com/t/mAWmT5bH

OpenStudy (anonymous):

You sure the B is -5i+2j-k instead of -5i+2j-4k ?

OpenStudy (christos):

bro notice A x 4B = -24i-68j-16k notice the 4B..

OpenStudy (anonymous):

well your teacher wrote down \[d _{1}\times 4d_{2}=(3i-2j+4k)\times(-20i+8j-16k)\] So I assume he wrote wrong.

OpenStudy (anonymous):

i would have to agree with zhantshen, clearly -16k should be -4k with the numbers that you gave